OpenStudy (anonymous):

How does sin(t)/(1-cos(t)) simplify to cos(t)/sin(t)?

OpenStudy (anonymous):

are you sure that's the exact question as written?

OpenStudy (anonymous):

It doesn't. Pick t=5 and you'll see it doesn't work.

OpenStudy (anonymous):

For example...

OpenStudy (anonymous):

Well it's that I'm looking at a solution my teacher wrote. He was showing how a cycloid had infinitely sharp cusps by finding the limit of the derivative of the parametric formulas given. So he was trying first to find the limit of sin(t)/(1-cos(t)) and he jumped to trying to find the limit of cos(t)/sin(t) because apparently, they're equal.

OpenStudy (anonymous):

Look at #5. http://people.sfcollege.edu/bruce.teague/pdf/2312%20W11/2312_MG5_solutions_W11.pdf This link should work.

OpenStudy (anonymous):

Did everybody leave?

OpenStudy (anonymous):

Awww c'mon!

OpenStudy (anonymous):

Now what am I gonna do?

OpenStudy (anonymous):

Calm down...I'll have a look.

OpenStudy (anonymous):

Haha ok

OpenStudy (anonymous):

Oh, he's used L'Hopital's rule...do you know about that?

OpenStudy (anonymous):

Maybe. I might have forgotten.

OpenStudy (anonymous):

I was about say that

OpenStudy (anonymous):

When you have indeterminate forms for the limit (numerator and denominator each go to something like 0/0 or infinity/infinity), the limit of the original ratio is equal to the limit ratio of the derivatives of the numerator and denominator.

OpenStudy (anonymous):

Take the derivative of the numerator of sin(t) and you get cos(t), and the derivative of 1-cost(t) is sin(t).

OpenStudy (anonymous):

Oh! It's all coming back now.

OpenStudy (anonymous):

OK. That makes sense. Thank you soooooo much.

OpenStudy (anonymous):

Fan me then! :) I want to reach superstar!

OpenStudy (anonymous):

How does one fan someone else?

OpenStudy (anonymous):

\[\lim_{t \rightarrow 0} \frac{sin(t)}{1-\cos(t)}= \frac{0}{1-1}=\frac{0}{0}\] lokisan is right, this is an indeterminate and using L'hopital would give you

OpenStudy (anonymous):

there should be a link-type thing next to my name..."Become a fan"

OpenStudy (anonymous):

I'm not seeing it. I'm looking.

OpenStudy (anonymous):

It's in the thread window...little 'thumbs up' icon. There's one next to your own name too.

OpenStudy (anonymous):

There. I got it.

OpenStudy (anonymous):

Awesome...cheers!